OpenStudy (lauluisa):

Which line is parallel to 6y +2x =-3

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (lauluisa):

\[A) 2y -\frac{ 1 }{ 3 }x-\frac{ 1 }{ 2 }\] \[B) 2y-6x=4\] \[C) 3y+x= -\frac{ 1 }{ 2 }\] \[D) y=3x+2\]

OpenStudy (anonymous):

C

OpenStudy (anonymous):

6y+2x=-3 Y-intercept form would be... y=-1/3x-1/2 C would be y=-1/3x-1/6 Parallel when they have the same slopes